1. Valles de Benasque y Eriste

OLD WAY FROM BENÀS (BENASQUE) TO SARLLÉ (CERLER) 1 hour 30 minutes

Take the track from the road to Los Baños, which is on the right of the road 50 m. before the Red Cross post. You will pass near the graveyard. Circumvent a knoll and the Old Benasque Castle that is on its right. Go up some small rocky edges and, after that, follow the old track that goes up to Cerler. After 25 min. cross the road to keep on ascending by the track. After approximately 7 more minutes take the road to the left and leave it again after 100 m. when you see another track on your right. After another 8 minutes take the main road again and follow it till you get to the town.